OpenStudy (anonymous):

Please help me multiply this rational expression: 8x/-1 * 1/4x^2

OpenStudy (anonymous):

is it really \[\huge \frac{8x}{-1}\times \frac{1}{4x^2}\]?

OpenStudy (anonymous):

usually you do not right "divided by -1" you would see instead \[-8x\times \frac{1}{4x^2}\] or \[\frac{-8x}{4x^2}\]

OpenStudy (anonymous):

cancelling gives you \[-\frac{2}{x}\]

OpenStudy (anonymous):

Okay great, could you help me with this one as well @satellite73 ? \[\frac{ 4s }{ 3t } x \frac{ 9s^2 }{ -4 }\]